# Scaling Laws for Uncertainty in Deep Learning

### v1

@@ 6 Conclusions
Inspired by recent works on scaling laws in deep learning, we investigated whether these apply to predictive uncertainties as well. Our empirical findings on computer vision and language tasks indicate that these scaling laws appear in a broad range of design choices, including architecture design, posterior approximation/ensembling approach, and hyper-parameters.
We set out to explain the emergence of such scaling laws and, inspired by slt, we provide novel theoretical insights on the connections between generalization error and predictive uncertainty. This helps us in understanding the power-law behavior associated with the scaling of information-theoretic measures of uncertainty with respect to dataset size.
While the theory gives some indications along these lines, the experiments reveal that obtaining exact power-law coefficients from the theory is not easy, as choices of model architecture, uq method, and hyper-parameters affect such coefficients in unpredictable ways. As a result, we believe that our work can be used to derive practical strategies to extrapolate uncertainties to large $N$, e.g., to predict how much data is needed to obtain ensemble predictions which are indistinguishable up to a given numerical precision.
In future works, we will investigate alternative theoretical angles to explain scaling laws on uncertainties in deep learning. One approach could be to study the behavior of uncertainties in models with large number of units, such as WideResNets, and connect their uncertainties with those obtained by Gaussian processes .
Limitations
Previous works on scaling laws for deep learning show clear trends of test performance with respect to model parameters; however, with our experimental setup, we did not observe the same for predictive uncertainties (see Fig. 7) and we were hoping to have more computational resources to investigate this aspect in more detail. Another element characterizing scaling laws for test performance is the behavior with respect to computational budget, which we do not consider here, again for reasons related to limited computational resources. Finally, it would have been great to study more applications where we could test the extrapolation of scaling laws to very large dataset sizes, in addition to the experiments on the GPT-2 language model.
Broader impacts.
This paper demonstrates the emergence of scaling laws of uncertainties in deep learning. The results should motivate a wider adoption of uq methods for deep learning, even in the case of supposedly large datasets, in order to obtain uncertainties that can be used to assist decision-making.
Appendix A Appendix
@@ A.1 Total Uncertainty in Bayesian Linear Regression
We consider a supervised learning problem with $N$ input-label training pairs $\{{{\boldsymbol{\mathbf{X}}}},{{\boldsymbol{\mathbf{y}}}}\}=\{({{\boldsymbol{\mathbf{x}}}}_{i},y_{i})\}_{i=1}^{N}$. Assuming that the data points are drawn independently from Eq. 15, the likelihood becomes:
@@
where ${{\boldsymbol{\mathbf{\mu}}}}_{N}$ and ${{\boldsymbol{\mathbf{\Sigma}}}}_{N}$ are the posterior mean and covariance matrix. From the predictive variance $\mathbb{V}[y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}}]$ we identify two components: (i) $\sigma^{2}$ which represents the uncertainty related to the data noise (au) and (ii) ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})^{\top}{{\boldsymbol{\mathbf{\Sigma}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})$ which is the uncertainty associated with model parameters (eu). As we discuss in Section 5.1, by defining $\mathbb{V}[y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}}]\coloneq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})$ as the predictive uncertainty for a new test point ${{\boldsymbol{\mathbf{x}}}}$ when training on $N$ input points, we have that $\sigma^{2}_{N+1}({{\boldsymbol{\mathbf{x}}}})\leq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})$, meaning that the posterior distribution becomes narrower as additional data points are observed (Qazaz et al., 1997). Moreover, $\lim_{N\to\infty}\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})=\sigma^{2}$; that is, in the infinite data regime, the predictive uncertainty converges to its irreducible component, au.
@@ A.2 Total Uncertainty and Generalization Error for linear models
[…]
We consider the same linear model of Section A.1 and without loss of generality we set the basis functions to be identity functions, i.e., ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})={{\boldsymbol{\mathbf{x}}}}$. We also introduce ${{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}}$, the set of underlying true parameters. A new data point $({{\boldsymbol{\mathbf{x}}}}_{N+1},y_{N+1})$ is generated according to the true process:
@@
[…]
@@
[…]
@@
This is closely related to the asymptotic expression we derived for tu in Section A.1.1 where, if we now consider identity basis functions and a test point ${{\boldsymbol{\mathbf{x}}}}_{N+1}$, it becomes:
@@ A.2.1 Asymptotic behaviors
In Watanabe (1999); Hayashi and Sawada (2025) it’s showed that the generalization error and the free energy have the asymptotic expansion:

### v2

@@ 6 Conclusions
Inspired by recent work on scaling laws in deep learning, we investigated whether similar patterns hold for predictive uncertainties. Across vision and language tasks, we find scaling behaviors robust to architecture, posterior approximation, and hyper-parameters. Building on slt, we provide theoretical insights connecting generalization error and predictive uncertainty, showing why information-theoretic measures scale with dataset size. However, deriving exact power-law coefficients is difficult, as they vary unpredictably with design choices. Our results suggest practical strategies for extrapolating uncertainties to large $N$, such as estimating how much data is needed for ensemble predictions to converge. They also point to applications in active learning, where one can predict the marginal uncertainty reduction from additional data to guide annotation budgets. More broadly, while scaling laws are predictive, they are not universal but depend on how the loss landscape is traversed, underscoring the need to tailor optimization strategies to each task.
Limitations and future work
A key limitation is that we do not yet have a formal theory that predicts the empirical scaling exponents $\gamma$ of deep networks from slt invariants or other geometric quantities. Moreover, optimization effects remain only partially characterized: we observe that simple optimization settings (e.g., fixed learning rates) can yield clearer scaling than more complex schedules in some settings, but we do not yet explain when and why this occurs. Finally, in transfer and fine-tuning settings, large-scale pre-training can saturate uncertainty estimates, potentially masking scaling behavior on smaller downstream datasets.
Future work will aim to (i) connect deep-network scaling exponents to effective model complexity (e.g., effective dimensionality or RLCT-like quantities), (ii) study more in depth the scaling with compute budget and model size, and (iii) relate uncertainty scaling to theoretical limits such as wide-network gp approximations.
Impact Statement
By characterizing uncertainty scaling in modern neural networks, this work provides essential insights for ensuring the reliability and safety of large-scale AI systems.
Appendix A Connections with slt
Our theoretical framework is motivated by slt, which offers geometric insights into the learning dynamics of highly overparameterized models. In deep neural networks, the Fisher information matrix is frequently singular; although these singularities form a measure-zero subset of the parameter space, they profoundly influence the learning process. slt demonstrates that the asymptotic properties of learning are governed by the algebraic geometry near these degenerate points (Watanabe, 2009, 2018).
In this section, we restrict our analysis to the connection between slt and tu within the context of Bayesian linear regression. Following the notation of Watanabe (2009), we define the nmll (or free energy) of a dataset ${{\boldsymbol{\mathbf{X}}}}_{N}$ of size $N$ as:
$F_{N}=-\log p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N})=-\log\int p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{\theta}}}})p({{\boldsymbol{\mathbf{\theta}}}})d{{\boldsymbol{\mathbf{\theta}}}}.$ (12)
The generalization error, $G_{N}$, measures the expected negative log-likelihood of an unseen data point $({{\boldsymbol{\mathbf{x}}}}_{N+1},y_{N+1})$ sampled from the true distribution. It is defined as:
$G_{N}$ $=\mathbb{E}_{p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{\theta}}}}_{\text{true}})}\left[-\log p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{y}}}}_{N})\right]$ (13)
$=-\mathbb{E}_{p(y)}\left[\log q_{N}(y)\right],$ (14)
where $p(y)$ denotes the true process $y_{N+1}\sim\mathcal{N}(y_{N+1}|{{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}}^{\top}{{\boldsymbol{\mathbf{x}}}}_{N+1},\sigma_{\mathrm{true}}^{2})$ and $q_{N}(y)$ represents the posterior predictive distribution given the first $N$ data points.
We can decompose the generalization error in Equation˜13 into aleatoric and epistemic components:
$G_{N}$ $=\mathbb{E}_{p(y)}[-\log{q_{N}(y)}]$
$=\mathbb{E}_{p(y)}[-\log p(y)+\log p(y)-\log{q_{N}(y)}]$
$=\underbrace{\mathbb{H}[p(y)]}_{\text{aleatoric uncertainty}}+\underbrace{\mathrm{KL}\big[p(y)||q_{N}(y)\big]}_{\text{epistemic uncertainty}}.$ (15)
Here, $\mathbb{H}[p(y)]=\frac{1}{2}\log{(2\pi e\sigma_{\mathrm{true}}^{2})}$ represents the irreducible noise (entropy) of the true process, while the kl divergence quantifies the eu arising from the deviation of the posterior predictive from the truth.
Examining the asymptotic behavior of tu as $N\rightarrow\infty$, and assuming identity basis functions without loss of generality, we obtain:
$\mathrm{TU}({{\boldsymbol{\mathbf{x}}}}_{N+1})=\mathbb{H}[q_{N}(y)]=\frac{1}{2}\log(2\pi e\sigma_{\mathrm{true}}^{2})+\frac{{{\boldsymbol{\mathbf{x}}}}_{N+1}^{\top}{{\boldsymbol{\mathbf{\Sigma}}}}_{X_{N+1}}^{-1}{{\boldsymbol{\mathbf{x}}}}_{N+1}}{2(N+1)}+O\left(\frac{1}{N^{2}}\right).$ (16)
As $N\to\infty$, the posterior predictive $q_{N}(y)$ converges to the true distribution $p(y)$. Consequently, the epistemic term vanishes, and both $G_{N}$ and tu asymptotically approach the irreducible aleatoric entropy. Watanabe (1999) derives a rigorous asymptotic expansion of $G_{N}$, showing it is proportional to the effective dimensionality of the model (see Section˜A.2.1).
Recent advances in estimating the effective dimensionality of deep models (Lau et al., 2024; Chen et al., 2024) appear especially promising. We intend to investigate these formal connections in future work, hypothesizing that concepts from Statistical Physics may further elucidate the observed scaling behaviors.
@@ A.1 Total Uncertainty in Bayesian Linear Regression
We consider a supervised learning problem with $N$ input-label training pairs $\{{{\boldsymbol{\mathbf{X}}}},{{\boldsymbol{\mathbf{y}}}}\}=\{({{\boldsymbol{\mathbf{x}}}}_{i},y_{i})\}_{i=1}^{N}$. Assuming that the data points are drawn independently from Equation˜19, the likelihood becomes:
@@
where ${{\boldsymbol{\mathbf{\mu}}}}_{N}$ and ${{\boldsymbol{\mathbf{\Sigma}}}}_{N}$ are the posterior mean and covariance matrix. From the predictive variance $\mathbb{V}[y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}}]$ we identify two components: (i) $\sigma^{2}$ which represents the uncertainty related to the data noise (au) and (ii) ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})^{\top}{{\boldsymbol{\mathbf{\Sigma}}}}_{N}{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{*})$ which is the uncertainty associated with model parameters (eu). As we discuss in Section˜5.1, by defining $\mathbb{V}[y_{*}|{{\boldsymbol{\mathbf{x}}}}_{*},{{\boldsymbol{\mathbf{y}}}}]\coloneq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})$ as the predictive uncertainty for a new test point $\boldsymbol{\mathbf{x}}$}} when training on $N$ input points, we have that $\sigma^{2}_{N+1}({{\boldsymbol{\mathbf{x}}}})\leq\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})$, meaning that the posterior distribution becomes narrower as additional data points are observed (Qazaz et al., 1997). Moreover, $\lim_{N\to\infty}\sigma^{2}_{N}({{\boldsymbol{\mathbf{x}}}})=\sigma^{2}$; that is, in the infinite data regime, the predictive uncertainty converges to its irreducible component, au.
@@ A.2 Total Uncertainty and Generalization Error for linear models
$F_{N}$ $=-\log p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N})=-\log\int p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N},\boldsymbol{\theta})p(\boldsymbol{\theta})d\boldsymbol{\theta}.$ (28)
We consider the same linear model of Section˜A.1 and without loss of generality we set the basis functions to be identity functions, i.e., ${{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}})={{\boldsymbol{\mathbf{x}}}}$. We also introduce ${{\boldsymbol{\mathbf{\theta}}}}_{\mathrm{true}}$, the set of underlying true parameters. A new data point $({{\boldsymbol{\mathbf{x}}}}_{N+1},y_{N+1})$ is generated according to the true process:
@@
$=-\log{\int{p({{\boldsymbol{\mathbf{y}}}}_{N}|{{\boldsymbol{\mathbf{X}}}}_{N},{{\boldsymbol{\mathbf{\theta}}}})p({{\boldsymbol{\mathbf{\theta}}}})d{{\boldsymbol{\mathbf{\theta}}}}}}$
$=-\log{\int{\left(\prod_{n=1}^{N}p(y_{n}|{{\boldsymbol{\mathbf{x}}}}_{n},{{\boldsymbol{\mathbf{\theta}}}})\right)p({{\boldsymbol{\mathbf{\theta}}}})d{{\boldsymbol{\mathbf{\theta}}}}}}.$ (30)
@@
$p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},D_{N})$ $=\int{p(y_{N+1}|{{\boldsymbol{\mathbf{x}}}}_{N+1},{{\boldsymbol{\mathbf{\theta}}}})p({{\boldsymbol{\mathbf{\theta}}}}|D_{N})d{{\boldsymbol{\mathbf{\theta}}}}}$
@@
This is closely related to the asymptotic expression we derived for tu in Section˜A.1.1 where, if we now consider identity basis functions and a test point ${{\boldsymbol{\mathbf{x}}}}_{N+1}$, it becomes:
@@ A.2.1 Asymptotic behaviors
In S. Watanabe (1999); 5 it’s showed that the generalization error and the free energy have the following asymptotic expansions: